# Stationary solutions to stochastic 3D Euler equations in Hölder space

### v1

@@ Abstract.
We establish the existence of infinitely many global and stationary solutions in $C(\mathbb{R};C^{\vartheta})$ space for some $\vartheta>0$ to the three dimensional Euler equations driven by an additive noise. The result is based on a new stochastic version of the convex integration method, incorporating the stochastic convex integration method developed in [HZZ22b] and pathwise estimates to derive uniform moment estimates independent of time.
@@ 1. Introduction
In this paper, we are considered with stochastic Euler equations governing the time evolution of the velocity $u$ of an inviscid fluid on the three dimensional torus $\mathbb{T}^{3}=\mathbb{R}^{3}/(2\pi\mathbb{Z})^{3}$. The system reads as
@@
where $P$ stands for the corresponding pressure. The right hand side represents a random external force acting on the fluid, which is given by a $GG^{*}$-Wiener process $B$ defined on some probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $GG^{*}$ being trace class operator. In early works, stochastic Euler equations have been widely studied in both two and three dimensions, see e.g.[BFM16], [BP01], [CFH19], [GHV14], where local-in-time solutions for (1.1) have been constructed. Our goal is to investigate the existence of global and stationary solutions with Hölder continuity to the above system (1.1). To achieve this, we employ the method of convex integration, a technique that has gained prominence in fluid dynamics over the past decades. This method was initially introduced following [DLS10, DLS12, DLS13] by De Lellis and Székelyhidi Jr. In these works, the method of convex integration was used in order to prove non-uniqueness of weak solutions with Hölder continuity to Euler equations. Subsequently, this method has led to several breakthroughs in the fluid dynamics, including proof of Onsager’s conjecture for the incompressible Euler equations [Ise18, BDLSV19]. We refer to the excellent review article by Buckmaster and Vicol [BV19] for a gentle introduction and further references. In recent years, convex integration has already been applied in a stochastic setting, namely, to the isentropic Euler system by Breit, Feireisl and Hofmanová [BFH20] and to the full Euler system by Chiodaroli, Feireisl and Flandoli [CFF19] with linear multiplicative noise. Futhermore, ill/well-posedness of dissipative martingale solutions to stochastic 3D Euler equations was studied by Hofmanová, Zhu and Zhu [HZZ22a], along with the existence and non-uniqueness of strong Markov solutions. Later, Hölder continuous, global-in-time probabilistically strong solutions to 3D Euler equations perturbed by Stratonovich transport noise was constructed in [HLP22]. In the mostly recent work [HZZ22b], existence and non-uniqueness of stationary solutions in $L^{2}$ to the stochastic 3D Euler equations was established, using a new stochastic convex integration method. However, the regularity of the global and stationary solutions to the stochastic 3D Euler equations with additive noise, as constructed in the previous work, cannot achieve Hölder continuity. Inspired by these previous works and the progress in the deterministic case, it is natural to investigate the existence and non-uniqueness of global and stationary Hölder continuous solutions to the Euler equations in the stochastic setting (1.1). Another motivation for our research is linked to the Onsager’s conjecture. It is constructed in [Ise18, BDLSV19] that solutions in $C^{\frac{1}{3}-}$ do not conserve energy to the deterministic Euler equations. It is natural to ask whether the regularity of the global, nonunique solutions in the stochastic setting can stay $C^{\frac{1}{3}-}$ a.s.. To summarize the above discussion, it is natural to investigate the validity of the following claims.
@@ 1.1. Main results
Our aim is to provide some answers to the above problems (i) and (ii) in the physically relevant context of the stochastic Euler equations on $\mathbb{T}^{3}$ driven by additive stochastic noise. As the main tool within this study, we apply a stochastic version of the convex integration method. Unlike the previous works using convex integration for the Navier–Stokes and Euler equations in the stochastic setting [HZZ19, HZZ22a, HZZ23a, HZZ23b], we are now inspired by [CDZ22] and [HZZ22b] where they overcome the limitation originating from stopping times, previously used to control the noise terms in the iteration. More precisely, we no longer work with stopping times and instead we include expectations in the inductive iteration and provide uniform moment estimates. However, the construction of Hölder continuous Euler flows in the deteministic setting, as demonstrated in [BV19], requires to introduce transport equations, and pathwise positive lower bounds for the solutions are used to improve regularity. As a result, we also introduce some pathwise estimates during iteration for convex integration. To this end, we use cutoff method to control the growth of noise pathwisely. Compared to only using stochastic convex integration, the combination of pathwise estimates with it can provide better Hölder regularity (see Remark 1.3 below). This approach also allows us not to control all the moments bounds during the iteration. We hope this cutoff method can be applied to enhance regularity of the solutions in the future work. As we are interested in the long time behavior of solutions and particularly in the construction of statistically stationary solutions, we work with entire solutions solving the equations for all times $t\in{\mathbb{R}}$. Accordingly, the norms in the convex integration scheme must be chosen appropriately in order to provide the desired global-in-time estimates. This is achieved through bounds of the form
$\sup_{t\in\mathbb{R}}\mathbf{E}\left[\sup_{t\leqslant s\leqslant t+1}\|u(s)\|_{C^{\vartheta}}^{p}\right]<\infty,$
with $p\geqslant 1$ and small $\vartheta>0$. Such bounds provide uniform moment estimates locally in $C(\mathbb{R};C^{\vartheta})$ and guarantee the convergence of the corresponding ergodic averages. This leads to the existence of stationary solutions. Within this study, we focus on analytically weak solutions which satisfy the equations in the following sense.
Definition 1.1.
We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is an analytically weak solution to the Euler system (1.1) provided
  1. (1)
  2. (2)
    $B$ is an $\mathbb{R}^{3}$-valued, spatial mean and divergence free, two-sided $GG^{*}$-Wiener process with respect to the filtration $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$;
  3. (3)
  4. (4)
    for every $-\infty<s\leqslant t<\infty$ it holds $\mathbf{P}$-a.s.
    $\langle u(t),\psi\rangle+\int_{s}^{t}\langle\mathord{{\rm div}}(u\otimes u),\psi\rangle{\mathord{{\rm d}}}r\ =\langle u(s),\psi\rangle+\langle B(t)-B(s),\psi\rangle$
    for all $\psi\in C^{\infty}(\mathbb{T}^{3})$, $\mathord{{\rm div}}\psi=0$.
With this definition at hand, our first main result reads as follows, tackling the problem (i) above.
@@
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exists an analytically weak solution to (1.1) in the sense of Definition 1.1 which belongs to $C(\mathbb{R},C^{\vartheta})$, for some $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}\right)$.
@@
(i) In Section 3, we prove this result by combining stochastic convex integration method with pathwise estimates to derive uniform moment estimates independent of time, and obtain Hölder continuous solutions, where $\vartheta$ can be chosen less than ${\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}$. While we also prove this result in a separable article (see [Lü24]), which applies stochastic convex integration method to the the construction of Hölder continuous Euler flows in [DLS13] to derive uniform moment estimates independent of time. The value of $\vartheta$ in this case decreases due to the higher moments required, and we may take $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{73{\cdot}62^{3}},\frac{1}{7{\cdot}62^{3}}\}}$. (ii) It is natural to ask whether we could construct the solutions with better regulartiy to (1.1) by combining the gluing method in [BDLSV19] with our cutoff method. This will be studied in our future work.
The second result is the existence and non-uniqueness of stationary solutions in Hölder space. In the sequel, we understand stationarity in the sense of shift invariance of laws of solutions on the space of trajectories, see also [BFHM19, BFH20e, FFH21, HZZ22b]. More precisely, we define the joint trajectory space for the solution and the driving Brownian motion as
${\mathcal{T}}=C({\mathbb{R}};C^{\kappa})\times C({\mathbb{R}};C^{\kappa})$
for some $\kappa>0$ and let $S_{t}$, $t\in{\mathbb{R}}$, be shifts on trajectories given by
@@
We note that the shift in the second component acts differently in order to guarantee that for a Brownian motion $B$ the shift $S_{t}B$ is again a Brownian motion.
@@
With these definitions at hand, we are able to state our second main result, tackling the problem (ii) above. The result then reads as follows and is proved in Section 5.
@@
In Section 2, we collect the basic notations used throughout the paper. Section 3 and Section 4 are the core of our proofs: the stochastic convex integration is developed and employed to construct entire non-unique analytically weak solutions. In Section 3, we present the main iteration Proposition 3.2 to prove our main Theorem 3.3. Section 4 is devoted to the construction of stochastic convex integration with pathwise estimates, namely, proof of iteration Proposition 3.2. Then the results in Section 3 are used in Section 5 together with a Krylov–Bogoliubov’s argument to obtain existence of non-unique stationary solutions to the stochastic Euler equations. In Appendix A, we recall the construction of Beltrami waves from [BV19] . In Appendix B, we give estimates on amplitude functions used in the convex integration construction. Finally, the estimates for transport equations are completed in Appendix C.
2. Notations
@@ 2.1. Function spaces
Throughout the paper, we employ the notation $a\lesssim b$ if there exists a constant $c>0$ such that $a\leqslant cb$. We let ${\mathbb{N}}_{0}:={\mathbb{N}}\cup\{0\}$. Given a Banach space $E$ with a norm $\|\cdot\|_{E}$ and $t\in{\mathbb{R}}$, we write $C_{t}E=C([t,t+1];E)$ for the space of continuous functions from $[t,t+1]$ to $E$, equipped with the supremum norm $\|f\|_{C_{t}E}=\sup_{s\in[t,t+1]}\|f(s)\|_{E}$. For $\alpha\in(0,1)$ we define $C^{\alpha}_{t}E$ as the space of $\alpha$-Hölder continuous functions from $[t,t+1]$ to $E$, endowed with the norm $\|f\|_{C^{\alpha}_{t}E}=\sup_{s,r\in[t,t+1],s\neq r}\frac{\|f(s)-f(r)\|_{E}}{|r-s|^{\alpha}}+\sup_{s\in[t,t+1]}\|f(s)\|_{E}$. For $\alpha\in(0,1)$, we also write $C^{\alpha}$ for functions in the space of $\alpha$-Hölder continuous functions from $\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$ endowed with the norm $\|f\|_{C^{\alpha}}=\sup_{x,y\in\mathbb{T}^{3},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in\mathbb{T}^{3}}|f(x)|.$ We use $L^{p}$ to denote the set of standard $L^{p}$-integrable functions from $\mathbb{T}^{3}$ to $\mathbb{R}^{3}$. For $s>0$, $p>1$ we set $W^{s,p}:=\{f\in L^{p};\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}<\infty\}$ with the norm $\|f\|_{W^{s,p}}=\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}$. Set $L^{2}_{\sigma}=\{f\in L^{2};\int_{\mathbb{T}^{3}}f\,{\mathord{{\rm d}}}x=0,\mathord{{\rm div}}f=0\}$. For $s>0$, we define $H^{s}:=W^{s,2}\cap L^{2}_{\sigma}$. For $s<0$ we define $H^{s}$ to be the dual space of $H^{-s}$. For $t\in\mathbb{R}$ and a domain $D\subset\mathbb{R}$ we denote by $C^{N}_{t,x}$ and $C^{N}_{D,x}$, respectively, the space of $C^{N}$-functions on $[t,t+1]\times\mathbb{T}^{3}$ and on $D\times\mathbb{T}^{3}$, respectively, $N\in\mathbb{N}_{0}$. The spaces are equipped with the norms
@@
For a Polish space $H$ we denote by $\mathcal{B}(H)$ the $\sigma$-algebra of Borel sets in $H$. We also use $\mathring{\otimes}$ to denote the trace-free part of the tensor product. We recall the inverse divergence operator $\mathcal{R}$ from [DLS13, Section 4.5], which acts on vector fields $v$ with $\int_{\mathbb{T}^{3}}v{\mathord{{\rm d}}}x=0$ as
@@
for $k,l\in\{1,2,3\}$. Then $\mathcal{R}v(x)$ is a symmetric trace-free matrix for each $x\in\mathbb{T}^{3}$, and $\mathcal{R}$ is a right inverse of the div operator, i.e. $\mathord{{\rm div}}(\mathcal{R}v)=v$. By [CL22, Theorem B.3] we have for $1\leqslant p\leqslant\infty$

### v3

@@ Abstract.
We establish the existence of infinitely many global and stationary solutions in $C(\mathbb{R},C^{\vartheta})$ space for some $\vartheta>0$ to the three-dimensional Euler equations driven by an additive stochastic forcing. The result is based on a new stochastic version of the convex integration method, incorporating the stochastic convex integration method developed in [HZZ22b] and pathwise estimates to derive uniform moment estimates independent of time.
@@ 1. Introduction
In this paper, we consider the stochastic Euler equations governing the time evolution of the velocity $u$ of an inviscid fluid on the three-dimensional torus $\mathbb{T}^{3}=\mathbb{R}^{3}/(2\pi\mathbb{Z})^{3}$. The system reads as
@@
where $P$ stands for the corresponding pressure. The right hand side represents a random external force acting on the fluid, which is given by a $GG^{*}$-Wiener process $B$ defined on some probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P})$ with $GG^{*}$ being trace class operator. The notion of solution to (1.1) we use throughout this paper is that of analytically weak solutions, here recalled.
Definition 1.1.
We say that $((\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in{\mathbb{R}}},\mathbf{P}),u,B)$ is an analytically weak solution to the stochastic Euler system (1.1) provided
  1. (1)
  2. (2)
    $B$ is an $\mathbb{R}^{3}$-valued, mean zero and divergence-free, two-sided $GG^{*}$-Wiener process with respect to the filtration $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$;
  3. (3)
  4. (4)
    for every $-\infty<s\leqslant t<\infty$ it holds $\mathbf{P}$-a.s.
    $\langle u(t),\psi\rangle=\langle u(s),\psi\rangle+\int_{s}^{t}\langle u,u\cdot\nabla\psi\rangle{\mathord{{\rm d}}}r+\langle B(t)-B(s),\psi\rangle$
    for all $\psi\in C^{\infty}(\mathbb{T}^{3},{\mathbb{R}}^{3})$, $\mathord{{\rm div}}\psi=0$.
@@
Stochastic Euler equations have been extensively studied in both two and three dimensions in previous works, including [BFM16, BP01, CFH19, GHV14], where local-in-time solutions for (1.1) have been constructed. Our objective is to investigate the existence of global and stationary solutions in the sense of Definition 1.1 with Hölder continuity to the above system (1.1). To achieve this goal, we employ the method of convex integration, a technique that has gained prominence in fluid dynamics over the past decades. This method was initially introduced following [DLS10, DLS12, DLS13] by De Lellis and Székelyhidi Jr. In these works, the convex integration method was used to prove the non-uniqueness of weak solutions with Hölder continuity to Euler equations. Subsequently, this method has led to several breakthroughs in fluid dynamics, including proof of Onsager’s conjecture for the incompressible Euler equations [Ise18, BDLSV19]. We refer to the excellent review article by Buckmaster and Vicol [BV19] for a gentle introduction and further references.
In recent years, convex integration has already been applied in a stochastic setting, namely, to the isentropic Euler system by Breit, Feireisl and Hofmanová [BFH20] and to the full Euler system by Chiodaroli, Feireisl and Flandoli [CFF21] with linear multiplicative noise. Furthermore, the ill/well-posedness of dissipative martingale solutions to stochastic 3D Euler equations was studied by Hofmanová, Zhu and Zhu [HZZ22a], along with the existence and non-uniqueness of strong Markov solutions. Later, Hölder continuous, global-in-time probabilistically strong solutions to 3D Euler equations perturbed by Stratonovich transport noise was constructed in [HLP24]. In the most recent work [HZZ22b], the existence and non-uniqueness of stationary solutions in $H^{\vartheta}$ to the stochastic 3D Euler equations was established, using a new stochastic convex integration method.
However, the regularity of the global and stationary solutions to the stochastic 3D Euler equations with additive noise, as constructed in the previous work, cannot achieve Hölder continuity. Inspired by these previous works and the progress in the deterministic case, it is natural to investigate the existence and non-uniqueness of global and stationary Hölder continuous solutions to the Euler equations in the stochastic setting (1.1). Another motivation for our research is linked to the Onsager’s conjecture. It is constructed in [Ise18, BDLSV19] that solutions in Hölder space $C^{\frac{1}{3}-}$ to the deterministic Euler equations do not conserve energy. It is natural to ask whether the global and nonunique solutions in the stochastic setting can stay $C^{\frac{1}{3}-}$ almost surely.
To summarize the above discussion, we investigate the validity of the following claims.
@@
Our objective is to address the aforementioned issues (i) and (ii) within the context of the stochastic Euler equations on $\mathbb{T}^{3}$ driven by additive stochastic noise. We employ a stochastic version of the convex integration method as the primary tool in this study. Unlike previous works applying convex integration for the Navier–Stokes and Euler equations in the stochastic setting [HZZ22a, HZZ23a, HZZ23b, HZZ24], our approach is inspired by [CDZ24] and [HZZ22b]. In these works, the limitations stemming from stopping times, which were previously used to control the noise terms in the iteration, are overcome. To be more specific, we have moved away from working with stopping times. Instead, we incorporate expectations into the inductive iteration and provide uniform moment estimates. However, the construction of Hölder continuous Euler flows in the deterministic setting, as demonstrated in [BV19], requires to introduce transport equations, and pathwise positive lower bounds for the solutions are used to improve regularity. As a result, we also introduce some pathwise estimates during iteration for convex integration and use the cutoff technique to control the growth of noise pathwisely. Compared to only using stochastic convex integration, the combination of pathwise estimates with it can provide better Hölder regularity of the solutions (see Remark 1.4 below). This approach also allows us not to control all the moment bounds during the iteration. We hope this cutoff method can be applied to enhance the regularity of the solutions in future work.
@@ 1.1. Main results
Our first main result reads as follows and is proved in Theorem 3.4, tackling the problem (i) above.
@@
Suppose that $\mathrm{Tr}((-\Delta)^{\frac{3}{2}+\sigma}GG^{*})<\infty$ for some $\sigma>0$, then there exists an analytically weak solution to (1.1) in the sense of Definition 1.1 with $(\mathcal{F}_{t})_{t\in{\mathbb{R}}}$ being the normal filtration generated by the Wiener process $B$, which belongs to $C(\mathbb{R},C^{\vartheta})$ for some $\vartheta\in\left(0,{\mathord{{\rm min}}}{\{\frac{\sigma}{120{\cdot}7^{5}},\frac{1}{3{\cdot}7^{5}}\}}\right)$.There are infinitely many such solutions $u$.
@@
In [HZZ22b], the authors have constructed solutions to (1.1) with continuous paths in a suitable Sobolev space, namely $C({\mathbb{R}},H^{\vartheta})$ $\mathbf{P}$-almost surely. In contrast, we are able to construct solutions achieving the Hölder continuity. Additionally, the stochastic convex integration developed in [HZZ22b] requires estimating all the moments of the solutions, while our analysis only necessitates specifying an iteration procedure for a finite number of moments, along with additional pathwise bounds.
Remark 1.4.
We can also establish Theorem 1.2 in a separate article (see [Lü24]), which sorely employs the stochastic convex integration method to provide uniform moment bounds. In other words, we do not incorporate pathwise estimates (see (3.8a), (3.8c), (3.8d) below) during the iteration for convex integration. In that paper, the construction of building blocks of the convex integration method follows from [DLS13], and we do not require pathwise lower bounds of the solutions to the transport equation (4.3) below. Therefore, the stochastic convolution $z$ is not cut off. However, the value of $\vartheta$ in this case decreases due to the arbitrarily higher moments required to be estimated, and we may take $\vartheta<{\mathord{{\rm min}}}{\{\frac{\sigma}{73{\cdot}62^{3}},\frac{1}{7{\cdot}62^{3}}\}}$.
Our second result establishes the existence and non-uniqueness of stationary solutions in the Hölder space. Here, stationarity is interpreted in terms of the shift invariance of laws governing solutions on the space of trajectories, as discussed in [BFHM19, BFH24, FFH21, HZZ22b]. To be more specific, we define the joint trajectory space for the solution and the driving Wiener process as follows:
${\mathcal{T}}=C({\mathbb{R}},C^{\kappa})\times C({\mathbb{R}},C^{\kappa})$
for some $\kappa>0$. And let $S_{t}$, $t\in{\mathbb{R}}$, be shifts on trajectories given by
@@
We observe that the shift in the second component operates differently to ensure that for a Wiener process $B$, the shift $S_{t}B$ remains a Wiener process.
@@
With these definitions at hand, our second main result then reads as follows and is proved in Theorem 6.1, tackling the problem (ii) above.
@@
Section 2 compiles the fundamental notations used throughout the paper. Section 3, 4, 5 constitute the core of our proofs, where stochastic convex integration is developed and employed to construct entire analytically weak solutions. In Section 3, we present the main iteration Proposition 3.3 to prove our main Theorem 3.4. Section 4 and Section 5 are devoted to the construction of stochastic convex integration with pathwise estimates, namely, proof of iteration Proposition 3.3. Then the results from Section 3 are applied in Section 6, coupled with a Krylov–Bogoliubov’s argument, to establish the existence of non-unique stationary solutions to the stochastic Euler equations. In Appendix A, we recall the construction and property of Beltrami waves from [BV19] . In Appendix B, we estimate amplitude functions used in the convex integration construction. Finally, the estimates for transport equations are completed in Appendix C.
2. Preliminaries
In the sequel, we use the notation $a\lesssim b$ if there exists a constant $c>0$ such that $a\leqslant cb$.
@@ 2.1. Function spaces
Given a Banach space $X$ with a norm $\|\cdot\|_{X}$ and $t\in{\mathbb{R}}$, we denote $C_{t}X:=C([t,t+1],X)$ as the space of continuous functions from $[t,t+1]$ to $X$, equipped with the supremum norm
$\|f\|_{C_{t}X}:=\sup_{s\in[t,t+1]}\|f(s)\|_{X}.$
For $\alpha\in(0,1)$, we use $C^{\alpha}_{t}X$ to denote the space of $\alpha$-Hölder continuous functions from $[t,t+1]$ to $X$, endowed with the norm
$\|f\|_{C^{\alpha}_{t}X}:=\sup_{s,r\in[t,t+1],s\neq r}\frac{\|f(s)-f(r)\|_{X}}{|r-s|^{\alpha}}+\|f\|_{C_{t}X}.$
We also write $C^{0}:=C(\mathbb{T}^{3},{\mathbb{R}}^{3})$ equipped with the supremum norm $\|f\|_{C^{0}}:=\sup_{x\in\mathbb{T}^{3}}|f(x)|$. For $\alpha\in(0,1)$, we write $C^{\alpha}:=C^{\alpha}(\mathbb{T}^{3},{\mathbb{R}}^{3})$ as the space of $\alpha$-Hölder continuous functions from $\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$ endowed with the norm
$\|f\|_{C^{\alpha}}:=\sup_{x,y\in\mathbb{T}^{3},x\neq y}\frac{|f(x)-f(y)|}{|x-y|^{\alpha}}+\sup_{x\in\mathbb{T}^{3}}|f(x)|.$
We denote $L^{p}$ as the set of standard $L^{p}$-integrable functions from $\mathbb{T}^{3}$ to $\mathbb{R}^{3}$. For $s>0$, $p>1$ the Sobolev space $W^{s,p}:=\{f\in L^{p};\|f\|_{W^{s,p}}:=\|(I-\Delta)^{{s}/{2}}f\|_{L^{p}}<\infty\}$. We set $L^{2}_{\sigma}:=\{f\in L^{2};\int_{\mathbb{T}^{3}}f\,{\mathord{{\rm d}}}x=0,\mathord{{\rm div}}f=0\}$. For $s>0$, we also denote $H^{s}:=W^{s,2}\cap L^{2}_{\sigma}$. For $t\in\mathbb{R}$ and a domain $D\subset\mathbb{R}$ and $N\in\mathbb{N}_{0}:={\mathbb{N}}\cup\{0\}$, we denote the space of $C^{N}$-functions from $[t,t+1]\times\mathbb{T}^{3}$ and $D\times\mathbb{T}^{3}$ to ${\mathbb{R}}^{3}$, respectively, by $C^{N}_{t,x}$ and $C^{N}_{D,x}$. The spaces are equipped with the norms
@@
We also use $\mathring{\otimes}$ to denote the trace-free part of the tensor product. For a tensor $T$, we denote its traceless part by $\mathring{T}:=T-\frac{1}{3}\mathrm{Tr}(T)\rm{Id}$.
We recall [DLS13, Definition 4.2] the inverse divergence operator $\mathcal{R}$ which acts on vector fields $v$ with $\int_{\mathbb{T}^{3}}v{\mathord{{\rm d}}}x=0$ as
@@
for $k,l\in\{1,2,3\}$. The above inverse divergence operator has the property that $\mathcal{R}v(x)$ is a symmetric trace-free matrix for each $x\in\mathbb{T}^{3}$, and $\mathcal{R}$ is a right inverse of the div operator, i.e. $\mathord{{\rm div}}(\mathcal{R}v)=v$. By [CL22, Theorem B.3] we have for $1\leqslant p\leqslant\infty$